OpenStudy (anonymous):

If theta is an angle in standard position and its terminal side passes through the point (-4,2) find the exact value of cos theta, and please show work it will really help.

OpenStudy (anonymous):

take the ratio of the adjacent side to the hypotenuse, but first we need the hypotenuse

OpenStudy (anonymous):

\[\sqrt{4^2+2^2}=\sqrt{16+4}=\sqrt{20}=2\sqrt{5}\]

OpenStudy (anonymous):

therefore \[\cos(\theta)=\frac{-4}{2\sqrt{5}}=\frac{-2}{\sqrt{5}}\]

OpenStudy (anonymous):

tysm

OpenStudy (anonymous):

yw